OpenStudy (anonymous):

Let f(x) = -6x + 3 and g(x) = 5x + 4. Find f*g and state its domain.

OpenStudy (anonymous):

Let f(x) = -6x + 3 and g(x) = 5x + 4. Find f*g and state its domain. A) -18x2 - 39x + 20; all real numbers B) -18x2 - 39x + 20; all real numbers except x = 1 C) -30x2 - 9x + 12; all real numbers D) -30x2 - 9x + 12; all real numbers except x = 4